# A Spatiotemporal, Quasi-experimental Causal Inference Approach to Characterize the Effects of Global Plastic Waste Export and Burning on Air Quality Using Remotely Sensed Data

### v2

@@ 1 Introduction
A decade ago, Wiedinmyer et al. (2014) presented the first comprehensive, global estimates of air pollution (and greenhouse gas) emissions due to open waste burning, using an emissions factor approach based on guidelines from the IPCC (Intergovernmental Panel on Climate Change). This approach relies on rough country-level approximations for the amount of waste generated and open-burned, which were derived as a deterministic function of national waste generation rate and waste collection efficiency estimates. The estimated amount of open waste burned was then converted into waste burning-attributable emissions estimates through multiplying by a literature-informed emissions factor. This and other inventories have subsequently been paired with chemical transport modeling to estimate the impact of open burning on regional air quality and attributable mortality (Gordon et al., 2023). While these studies provide evidence that the air pollution contributions of waste burning may be substantial, their methods require strong assumptions backed up by very limited empirical data, and adaptation of these methods to conduct policy impact evaluations, which are crucial for future policy making, would necessitate further strong assumptions about the impact of a policy on the various inputs.
In this study, we focus on plastic waste, whose global generation more than doubled from 2000-2019 (OECD, 2022) and whose transport / trade from higher- to lower-income countries is a major environmental justice issue. In contrast to earlier studies estimating emissions from burning all kinds of waste, recent studies have focused on more robust estimation of plastic waste generation in specific regions (Bardales Cruz et al., 2023) and detecting molecular tracers from plastic burning (Islam et al., 2022). However, these methods do not easily scale to large regions and/or time periods.
@@
Motivated by this case study, we also propose a new, multiply-robust causal inference approach to estimation and inference for quasi-experimental study designs with a “universal” but dose-dependent intervention. A growing body of work uses quasi-experimental study designs to estimate the air quality impact of a policy or intervention, with a recent surge focused on the Covid-19 lockdown – which resembles our setting in causing a nationwide shift in emissions-relevant behaviors (Heffernan et al., 2024; Chen et al., 2021; Dey et al., 2021). Policies that are adopted “universally” across a country/region of interest present a challenge for many conventional quasi-experimental analytic methods, which typically assume the existence of a control group. To address this, Heffernan et al. (2024) developed a framework for machine learning-based comparative interrupted time series (CITS), in which U.S. city-specific models were used to estimate the difference in average daily air pollution before and after the lockdown dates, using past years of data as controls. A closely related approach was taken by Chen et al. (2021), who used synthetic controls in place of CITS. However, these methods that model each location’s effect separately do not satisfy our needs, because we are interested in characterizing the relationship between changes in air quality and each location’s “dose” of exposure to the treament / intervention, which in our case is port proximity as defined in Section 2.3. That is, we wish to estimate a causal exposure-response curve.
Our proposed method merges ideas from this literature on quasi-experiments with a “universal” intervention and ideas from the burgeoning literature on multiply-robust causal exposure-response curve estimation. Specifically, we build closely upon recent work by Hettinger et al. (2025), who introduced a multiply-robust estimator for causal exposure-response curves within the difference-in-differences (DiD) framework. Derived from the efficient influence function or EIF (see Section 3.2), this estimator accounts for confounding both between treated and control units and across levels of exposure to the treatment / intervention, which can be thought of as a dosage. We propose a related EIF-based multiply-robust estimator for causal exposure-response curves in quasi-experimental designs with a universal intervention. We also propose an uncertainty quantification approach that accounts for both spatial and temporal correlation. We apply this method to remotely sensed data for the Indonesia plastic waste case study to estimate the change in air pollution at open dump sites due to increased plastic waste in the wake of China’s 2018 ban, as a function of a dump site’s port proximity.
@@ 2.1.1 Why Non-dump Sites Should Not Be Used as “Control” Locations
To motivate our methodological approach (in which we use past years as controls for future years at the GPW dump sites), briefly consider what trying to identify separate “control” locations would look like.
Let us walk through several of the causal identifying assumptions made in Hettinger et al. (2025), which are standard in DiD-type analyses. Positivity requires that controls have a non-zero probability of being treated. In our setting, this might be reasonable if we selected non-dump locations with similar characteristics to dump sites, including their proximity to major roads and population density. The Stable Unit Treatment Value Assumption (SUTVA) requires that (i) there is no interference between units and (ii) there are no hidden variations in treatment (Rubin, 1980). In our setting, spillover of air pollution becomes a concern if controls are spatially proximate to dump sites, which would be made more likely if we prioritized non-dump locations with similar landcover characteristics. The second aspect of SUTVA might be violated in our case study if we accidentally selected some “control” locations where waste is dumped and burned, but dumping there went undetected in the satellite imagery. Conditional Parallel Trends requires that the pre-post difference in the treated group would be the same (conditional on covariates) as that in the control group, in the absence of treatment. Here, consider that Indonesia’s domestic plastic waste generation is increasing by about 5% per year (Global Plastic Action Partnership, 2020). Since we know that a substantial fraction of this waste is taken to dump sites and burned, we would expect the long-term trend in air quality over dump sites to be different than that over non-dump sites, regardless of waste being imported.
Taken together, these considerations illustrate why attempting to use non-dump sites as controls would likely lead to violations of causal identifying assumptions and biased estimates. Thus, we use only waste dump sites in our analysis and assume that they are universally treated by the China plastic waste ban starting in 2018.
@@
We extract these monthly PM2.5 values for 2012-2019 at the GPW sites, for each site taking the value of the grid cell in which the site centroid lies. The average PM2.5 concentration at the dump sites in 2012-2017 (prior to the China ban) was 25.7 $\mu g/m^{3}$. The time series of concentrations averaged across sites (Figure 1) shows a strong seasonal pattern across our study period.
Figure 1: Median, mean, and 75th percentile of PM2.5 across GPW sites in Indonesia, 2012-2019.
@@ 2.3 Exposure: Port Proximity
Because the policy intervention is on waste imports and in Indonesia any such imports should come through ports, presumably dump sites closer to port activity should be more heavily affected by waste imports. Thus, we consider each dump site’s proximity to port activity to be a proxy measure of treatment dose (or “exposure”). Following other recent work assessing the domestic impacts of China’s plastic waste ban Unfried and Wang (2022), we could use distance to the nearest port as a proxy. However, that would not take into account the amount of shipping activity in each location. Therefore, we used data from shipping signals remotely sensed via the Automated Identification System, specifically a derived measure of cargo ship loitering from the Global Maritime Traffic Density Service (2021). For our continuous exposure metric, we calculated an inverse-distance-weighted sum of the 2018 loitering data around every GPW site. Due to the strongly right-skewed distribution of this port proximity index, we use a quantile version throughout the analysis, as shown in Figure 2. This transformed variable is henceforth referred to simply as the “port proximity”.
@@ 2.4 Covariates: Meteorology, Population Density, and Fire Locations
To assess plausibility that changes in waste burning are contributing to any changes in air quality post-intervention, we conduct a post-hoc exploratory analysis of the temporal trend in the number of VIIRS active fire points at open dump sites in Indonesia during the study period. Increases in active fire points at open dump sites from 2018 onwards would provide informal evidence that changes in air quality may be (at least partially) due to increases in waste burning.
@@ 3.1 Setup and Assumptions
Let $\mathcal{O}$ be the observed data, containing: $Y$ the outcome (PM2.5), $D$ the dose of exposure (port proximity), and $\mathbf{X}$ the covariates. The latter includes the population living within 1km of the site, monthly averaged meteorology (temperature, humidity, precipitation, pressure, and wind speed), and the number of active fires detected in the province (that month) which did not overlap with any GPW site. To help capture spatial heterogeneity, we also include fixed effects for site ID and province in our model; the spatial boundaries for the latter are shown in Figure S3. Lastly, to account for longer-term trends, we include time $t=1,...,96$ to represent the months between 2012 and 2019, inclusive.
@@
Furthermore, let $B$ (for “ban”) be a pre- vs. post-intervention indicator. $B=0$ indicates dates before policy implementation, and $B=1$ indicates dates after (in our case, 1/1/2018 onwards). Because all units (dump sites) in the analysis are treated post-intervention, $B$ is also a binary indicator of treatment status. We use $B$ rather than the more commonly used $A$ to illustrate this departure from DiD.
@@
Each observation is indexed by its location $i$ ($N$ = 356 dump sites) and time point $t$. When $B_{it}=0$, we say $D_{it}=\emptyset$, meaning $D$ is undefined – while each location $i$ has a port proximity prior to China’s ban, this feature has not yet been activated to represent the dose of treatment. We define the potential outcomes as $Y_{it}(B_{it}=b,D_{it}=\delta)$, i.e., the outcome that would have been observed for a given unit under intervention status $b$ and dose $\delta$. This definition of the potential outcomes requires that the future cannot affect the past (arrow of time, call this A1) and SUTVA (A2), discussed in section 2.1.1. Mathematically, SUTVA requires that $Y_{it}(\mathbf{B}_{t},\mathbf{D}_{t})=Y_{it}(B_{it},D_{it})$, where $\mathbf{B}_{t}$ and $\mathbf{D}_{t}$ denote the population-level vectors of intervention and dosage.
@@
Our aim is to estimate a causal exposure-response curve quantifying the causal effect on Indonesian air quality of China’s 2018 plastic waste import ban at different levels of port proximity, $\delta$. In the DiD setting, Hettinger et al. (2025) terms this the “Average Dose Effect on the Treated”, or ADT. While noting again that our formulation is different, we still see fit to call our effect of interest the ADT:
$ADT(\delta)=\Psi(\delta)=E[Y(1,\delta)-Y(0,\emptyset)|B=1]$
@@
In our case study, this estimand describes the average difference in air quality post-China ban at open dump sites in Indonesia, if all dump sites were exposed to port proximity dose $\delta$. To identify the $ADT(\delta)$, we require
[…]
@@
[…]
This means that there is no residual time-varying confounding. Assembling an adequate set of covariates to justify this assumption relies on domain knowledge. Its plausibility can be strengthened by adjusting for seasonal and/or long-term trends, e.g., through inclusion of a time variable in the adjustment set. This assumption is needed due to the absence of control locations in our application.
@@
$E[Y(1,\delta)|B=1,D=\delta,\mathbf{X}]=E[Y(1,\delta)|B=1,\mathbf{X}]\>\>\forall\>\>\delta\in\mathbb{D}$
This means that there is no residual spatial confounding, in other words we would expect the post-policy air quality to be the same at all dump sites with covariates $\mathbf{X}$, if they were assigned dose $\delta$ (regardless of their observed port proximity). The plausibility of this assumption may be strengthened by adjusting for region-specific fixed effects (in our case, province).
@@
$\Psi(\delta)$ $=E[Y(1,\delta)-Y(0,\emptyset)|B=1]$ (1)
$=E[E[Y(1,\delta)-Y(0,\emptyset)|B=1,\mathbf{X}]|B=1]$ (2)
$=E[E[Y(1,\delta)|B=1,D=\delta,\mathbf{X}]-E[Y(0,\emptyset)|B=0,\mathbf{X}]|B=1]$ (3)
@@
Where line (1) is by definition, (2) by iterated expectations, (3) by A6 on the lefthand side and A5 on the righthand side, and (4) by A1-3.

### v3

@@ Abstract
Keywords: Air Pollution, Causal Inference, Plastic Waste, Policy Evaluation, Remote Sensing, Spatiotemporal
@@ 1 Introduction
A decade ago, Wiedinmyer et al. (2014) presented the first comprehensive, global estimates of air pollution (and greenhouse gas) emissions due to open waste burning, using an emissions factor approach based on guidelines from the IPCC (Intergovernmental Panel on Climate Change). This approach relies on rough country-level approximations for the amount of waste generated and open-burned, which were derived as a deterministic function of national waste generation rate and waste collection efficiency estimates. The estimated amount of open waste burned was then converted into waste burning-attributable emissions estimates through multiplying by a literature-informed emissions factor. This and other inventories have subsequently been paired with chemical transport modeling to estimate the impact of open burning on regional air quality and attributable mortality (Gordon et al., 2023). While these studies provide evidence that the air pollution contributions of waste burning may be substantial, their methods require strong assumptions backed up by very limited empirical data (e.g., that all countries around the world burn approximately 60% of the waste that is available to be burned), and adaptation of these methods to conduct policy impact evaluations, which are crucial for future policy making, would necessitate further strong assumptions about the impact of a policy on the various inputs.
In this study, we focus on plastic waste, whose global generation more than doubled from 2000-2019 (OECD, 2022) and whose transport / trade from higher- to lower-income countries is a major environmental justice issue. In contrast to earlier studies estimating emissions from burning all kinds of waste, recent studies have focused on more robust estimation of plastic waste generation in specific regions (Bardales Cruz et al., 2023) and detecting molecular tracers from plastic burning (Islam et al., 2022). However, these methods do not easily scale to large regions and/or time periods. For context, Bardales Cruz et al. (2023) surveyed 50 households in one rural indigenous community in Guatemala over 4 weeks and Islam et al. (2022) deployed an air quality monitor sequentially at two sites in Bangladesh, leaving it in place for 3-4 months at each site.
@@
Motivated by this case study, we also propose a new, multiply-robust causal inference approach to estimation and inference for quasi-experimental study designs with a “universal” but dose-dependent intervention. A growing body of work uses quasi-experimental study designs to estimate the air quality impact of a policy or intervention, with a recent surge focused on the Covid-19 lockdown – which resembles our setting in causing a nationwide shift in emissions-relevant behaviors (Heffernan et al., 2024; Chen et al., 2021; Dey et al., 2021). Policies that are adopted “universally” across a country/region of interest present a challenge for many conventional quasi-experimental analytic methods, which typically assume the existence of a control group. To address this, Heffernan et al. (2024) developed a framework for machine learning-based comparative interrupted time series (CITS), in which U.S. city-specific models were used to estimate the difference in average daily air pollution before and after the lockdown dates, using past years of data as controls. A closely related approach was taken by Chen et al. (2021), who used synthetic controls in place of CITS. Similarly, Ma et al. (2021) used regression discontinuity in time to estimate the air quality impact of a new public transport provision in London, at different sites. However, these methods that model each location’s effect separately do not satisfy our needs, because we are interested in characterizing the relationship between changes in air quality and each location’s “dose” of exposure to the treatment / intervention, which in our case is port proximity as defined in Section 2.3. That is, we wish to estimate a causal exposure-response curve.
Our proposed method merges ideas from this literature on quasi-experiments with a “universal” intervention and ideas from the burgeoning literature on multiply-robust causal exposure-response curve estimation. Specifically, we build closely upon recent work by Hettinger et al. (2025a), who introduced a multiply-robust estimator for causal exposure-response curves within the difference-in-differences (DiD) framework. Derived from the efficient influence function or EIF (see Section 3.2), this estimator accounts for confounding both between treated and control units and across levels of exposure to the treatment / intervention, which can be thought of as a dosage. We propose a related EIF-based multiply-robust estimator for causal exposure-response curves in quasi-experimental designs with a universal intervention. We also propose an uncertainty quantification approach that accounts for both spatial and temporal correlation. We apply this method to remotely sensed data for the Indonesia plastic waste case study to estimate the change in air pollution at open dump sites due to increased plastic waste in the wake of China’s 2018 ban, as a function of a dump site’s port proximity.
@@ 2.1.1 Why Non-dump Sites Should Not Be Used as “Control” Locations
To motivate our methodological approach (in which we use past years as controls for future years at the GPW dump sites), briefly consider what trying to identify separate “control” locations would look like. Let us walk through several of the causal identifying assumptions made in Hettinger et al. (2025a), which are standard in DiD-type analyses.
Positivity requires that controls have a non-zero probability of being treated, in other words, that non-dump sites selected as controls had some chance of receiving import-related plastic waste (which could either be imported waste or domestic waste displaced by imported waste) as a result of China’s ban. In practice, we would need to be able to select non-dump locations with appropriate overlap of key confounders (e.g., population density and climate characteristics) such that they could be considered “similar” locations to dump sites.
The Stable Unit Treatment Value Assumption (SUTVA) requires that (i) there is no interference between units and (ii) there are no hidden variations in treatment (Rubin, 1980). In our setting, spillover of air pollution becomes a concern if controls are spatially proximate to dump sites, which would be made more likely if we prioritized non-dump locations with similar landcover characteristics. The second aspect of SUTVA might be violated in our case study if we accidentally selected some “control” locations where waste is dumped and burned, but dumping there went undetected in the satellite imagery.
Conditional Parallel Trends requires that the pre-post difference in the treated group would be the same (conditional on covariates) as that in the control group, in the absence of treatment. Here, consider that Indonesia’s domestic plastic waste generation is increasing by about 5% per year (Global Plastic Action Partnership, 2020). Since we know that a substantial fraction of this waste is taken to dump sites and burned, we would expect the long-term trend in air quality over dump sites to be different than that over non-dump sites, regardless of waste being imported.
Taken together, these considerations illustrate why attempting to use non-dump sites as controls would likely lead to violations of causal identifying assumptions and biased estimates. Thus, we use only waste dump sites in our analysis and assume that they are universally treated by the China plastic waste ban starting in 2018. Note that we do not expect any dump sites to be sufficiently far away from ports to be unaffected by import-related plastic waste (such that they could be used as control locations) because Indonesia is an island nation and many plastic recycling facilities are known to be far inland (Septiono and Ismawati, 2022). Empirically, 98% of GPW sites are within 100km of the coast; the farthest site is 150km inland.
As an aside, it would be equally if not more problematic to try to use control locations in other countries because of differing (rates of change of) waste generation and burning practices, especially as many other countries in the East Asia & Pacific region were similarly exposed to the effects of China’s ban.
@@
We extract these monthly PM2.5 values for 2012-2019 at the GPW sites, for each site taking the value of the grid cell in which the site centroid lies. The average PM2.5 concentration at the dump sites in 2012-2017 (prior to the China ban) was 25.7 $\mu g/m^{3}$.
The time series of concentrations averaged across sites (Figure 1) shows a strong seasonal pattern across our study period. However, without accounting for other time-varying factors, we cannot tell from a plot like this whether there was an unexpected increase in PM2.5 in 2018-2019.
Figure 1: Median, mean, and 75th percentile of PM2.5 across GPW [dump] sites in Indonesia, 2012-2019.
@@ 2.3 Exposure: Port Proximity
Because the policy intervention is on waste imports and in Indonesia any such imports should come through ports, presumably dump sites closer to port activity should be more heavily affected by waste imports. Thus, we consider each dump site’s proximity to port activity (described below) to be a proxy measure of treatment dose or “exposure”. Note that to follow other recent work assessing the domestic impacts of China’s plastic waste ban (Unfried and Wang, 2022), we could use distance to the nearest port as a proxy; however, that would not take into account the amount of shipping activity in each location. Therefore, we used data from shipping signals remotely sensed via the Automated Identification System, specifically a derived measure of cargo ship loitering from the Global Maritime Traffic Density Service (2021). For our continuous exposure metric, we calculated an inverse-distance-weighted sum of the 2018 loitering data around every GPW site, up to a threshold of 100km. Due to the strongly right-skewed distribution of this port proximity index, we use a quantile version throughout the analysis, as shown in Figure 2. This transformed variable is henceforth referred to simply as the “port proximity”.
@@ 2.4 Covariates: Meteorology, Population Density, and Fire Locations
To assess plausibility that changes in waste burning are contributing to any changes in air quality post-intervention, we conduct a complementary exploratory analysis of the temporal trend in the number of VIIRS active fire points at open dump sites in Indonesia during the study period. Increases in active fire points at open dump sites from 2018 onwards would provide informal evidence that changes in air quality may be (at least partially) due to increases in waste burning.
@@ 3.1 Setup and Assumptions
Let $\mathcal{O}$ be the observed data, containing: $Y$ the outcome (PM2.5), $D$ the dose of exposure (port proximity), and $\mathbf{X}$ the covariates. The latter includes the population living within 1km of the site, monthly averaged meteorology (temperature, humidity, precipitation, pressure, and wind speed), and the number of active fires detected in the province (that month) which did not overlap with any GPW site. To help capture spatial heterogeneity, we also include fixed effects for site ID and province in our model; the spatial boundaries for the latter are shown in Figure S3. This is considered a reasonable approach in settings with many units where only few of these are important/informative (Fuhr and Papies, 2024) and when both bigger-scale (e.g., province level) and smaller-scale (individual dump site level) heterogeneity are likely at play. Both of these characteristics are present in our case study, as shown by a complementary exploratory analysis (described in Section 4.2) illustrating that the distribution of the number of large fires (detectable by VIIRS) occurring at each dump site is highly skewed, with some clustering by province. Lastly, to account for longer-term trends, we include time t=1,,96t=1,...,96 to represent the months between 2012 and 2019, inclusive.
@@
Furthermore, let $B$ (for “ban”) be a pre- vs. post-intervention indicator. $B=0$ indicates dates before policy implementation, and $B=1$ indicates dates after (in our case, 1/1/2018 onwards). Because all units (dump sites) in the analysis are treated post-intervention, $B$ is also a binary indicator of treatment status. We use $B$ rather than the more commonly used $A$ to illustrate this departure from DiD. As mentioned in the Introduction, while our use of time-varying covariates and pre-intervention years as controls for post-intervention years shares features with the synthetic control and interrupted time series literatures, existing methods do not enable estimating an exposure-response curve across locations.
@@
Each observation is indexed by its location $i$ (NN = 356 dump sites) and time point $t$. When $B_{it}=0$, we say $D_{it}=\emptyset$, meaning $D$ is undefined – while each location $i$ has a port proximity prior to China’s ban, this feature has not yet been activated to represent the dose of treatment. We define the potential outcomes as $Y_{it}(B_{it}=b,D_{it}=\delta)$, i.e., the outcome that would have been observed for a given unit under intervention status $b$ and dose $\delta$. This definition of the potential outcomes requires that the future cannot affect the past (arrow of time, call this A1) and SUTVA (A2), discussed in Section 2.1.1. Mathematically, SUTVA requires that $Y_{it}(\mathbf{B}_{t},\mathbf{D}_{t})=Y_{it}(B_{it},D_{it})$, where $\mathbf{B}_{t}$ and $\mathbf{D}_{t}$ denote the population-level vectors of intervention and dosage. Note that while spillover of air pollution between dump sites would violate A2, this is unlikely to substantially impact our final effect curve because sites close together have similar ”dose” values of port proximity, so our estimation procedure will average over them.
In order to avoid structural ill-definition, we also introduce an indicator for $t\geq T_{0}$, where $T_{0}$ is the treatment initiation time. In practice, $T_{0}$ is 1/1/2018 for all sites, but conceptually this could have occurred at any point in our study period, such that any time point could have been untreated.
@@
Our aim is to estimate a causal exposure-response curve quantifying the causal effect on Indonesian air quality of China’s 2018 plastic waste import ban at different levels of port proximity, $\delta$. In the DiD setting, Hettinger et al. (2025a) terms this the “Average Dose Effect on the Treated”, or ADT. While noting again that our formulation is different, we still see fit to call our effect of interest the ADT:
$ADT(\delta)=\Psi(\delta):=E[Y_{it}(1,\delta)-Y_{it}(0,\emptyset)|t\geq T_{0}]$
@@
In our case study, this estimand describes the average difference in air quality post-China ban at open dump sites in Indonesia, if all dump sites were exposed to port proximity dose $\delta$.
An important note about the ADT is how it differs from a more commonly-used Conditional Average Treatment Effect, $CATE(\delta):=E[Y_{it}(B=1)-Y_{it}(B=0)|D_{it}=\delta]$. First, recall that $D$ is undefined when $B=0$, not yet having been ‘activated’. Second, foreshadowing the following sections, the ADT approach enables interpreting the final effect curve in terms of port proximity versus in terms of port proximity-possibly-confounded-by-other-factors. It does so by estimating a generalized propensity score for $D$ and a binary propensity score for $B$, whereas a traditional doubly-robust estimator for the CATE would only use the latter.
In addition to A1 (arrow of time) and A2 (SUTVA), we assume
$\text{(A3) Consistency: }Y_{it}(b,\delta)=Y_{it}\text{ when }(B_{it},D_{it})=(b,\delta)$
$\text{(A4) Positivity: there exists }\epsilon>0\text{ such that}$
$\>\>\>\>\text{(i) }\epsilon\leq\pi_{B}(\mathbf{x})<1\>\forall\>\mathbf{x}\in\mathbb{X},\>\>\>\>\>\>\>\>\pi_{B}(\mathbf{x})=P(B=1|\mathbf{x})$
$\>\>\>\>\text{(ii) }\epsilon\leq\pi_{D}(\delta|\mathbf{x},B=1)\>\forall\>\mathbf{x}\in\mathbb{X}|B=1,\delta\in\mathbb{D},\>\>\>\>\>\>\>\>\pi_{D}(\delta|\mathbf{x},B=1)=p(D=\delta|B=1,\mathbf{X}=\mathbf{x})$
@@
$\text{(A5) Conditional Ignorability Between Time Periods: }Y_{it}(0,\emptyset)\perp\!\!\!\perp I(t\geq T_{0})\>|\>\mathbf{X}_{it}$
This means that there is no unmeasured time-varying confounding – no unmeasured factor causing air pollution to be systematically different in the periods before and after China’s ban, after accounting for confounding factors. Assembling an adequate set of covariates to justify this assumption relies on domain knowledge. Its plausibility can be strengthened by adjusting for seasonal and/or long-term trends, e.g., through inclusion of a time variable in the adjustment set. Note that in this setting, we do not have to worry about time-varying confounding affected by past exposure, because China’s ban being enacted should not affect confounders like meteorology and population density in Indonesia. However, we must pay careful attention to any other policies with the potential to make air pollution, conditional on covariates, systematically different in the $B=0$ and $B=1$ time periods – see discussion of this in Section 5.
@@
$Y_{it}(1,\delta)\perp\!\!\!\perp D_{it}\>|\>t\geq T_{0},\mathbf{X}_{it}\>\>\>\>\>\>\>\>\forall\>\>\delta\in\mathbb{D}$
This means that there is no unmeasured spatial confounding, in other words we would expect the post-policy air quality to be the same at all dump sites with covariates $\mathbf{X}$, if they were assigned dose $\delta$ (regardless of their observed port proximity). The plausibility of this assumption may be strengthened by adjusting for region-specific fixed effects (in our case, province).
Conditional ignorability (in both A5 and A6) is a stronger assumption than parallel trends, in a setting where control locations are available. However, for our application, the assumption that some locations represent the control condition post-treatment (China ban) would also be strong and likely inaccurate, as described in Section 2.1.1. Note that while we believe it is reasonable to assume conditional ignorability on distributions here given our domain knowledge, all that is needed for identifiability of our estimand is (a weaker) conditional ignorability in expectations, as is shown below.
@@
$\Psi(\delta)$ $=E[Y(1,\delta)-Y(0,\emptyset)|t\geq T_{0}]$ (1)
$=E[E[Y(1,\delta)-Y(0,\emptyset)|t\geq T_{0},\mathbf{X}]|t\geq T_{0}]$ (2)
$=E[E[Y(1,\delta)|t\geq T_{0},D=\delta,\mathbf{X}]-E[Y(0,\emptyset)|t<T_{0},\mathbf{X}]|t\geq T_{0}]$ (3)
@@
Where line (1) is by definition, (2) by iterated expectations, (3) by A6 on the lefthand side and A5 on the righthand side, and (4) by definition of $B$ and A1-3.
Finally, note that our port proximity index, $D$, is a proxy of the true continuous exposure variable of interest, i.e. the amount of China ban-instigated import-related plastic waste, which could either be imported waste or domestic waste displaced by imported waste. (Figure S5 illustrates this with a directed acyclic graph.) We denote this true exposure $G$. Thus, our use of $D_{it}$ as a proxy for $G_{it}$ requires
(A7) Valid Proxy for the Continuous Exposure: $D$ and $G$ satisfy both of the following
$\>\>\>\>\text{(i) }D_{it}\not\!\perp\!\!\!\perp G_{it}$
$\>\>\>\>\text{(ii) }Y_{it}\perp\!\!\!\perp D_{it}\>|\>G_{it},X_{it}$